ePESSTO Year 1+2 - Final SOFI Spectra reduction


Summary:
SOFI spectra reduced with PESSTO pipeline version 2.4.0.

The final SOFI reductions resulted in 83 reduced near-IR science spectra of which 62 in GB, and 21 in GR. Notes on the reductions below.


2017-05-17:
SN2017cbv GB and GR observations did not include accompanying telluric or spectroscpic standard star observations during the same night. Using Hip94200 and EG274 from 2017-04-23.

2017-08-19:
Feige110 spectroscpic standard star GB spectral extraction appears to have a potential artifact. EG274 2017-08-21 spectrum was used for the calibration of AT2017fzw and AT2017gah spectra from 2018-08-19.

2017-08-21:
PS17egl GB sequence of 16 frames is decreasing in quality. The last 4 frames were excluded from the reduction.

2017-08-28:
EG274 spectroscopic standard star GB spectral reduction fails. EG274 2017-08-21 spectrum was used for the calibration of AT2017gah spectrum from 2018-08-28.

2017-08-29:
Feige110 spectroscpic standard star GB spectral extraction appears to have a potential artifact. EG274 2017-08-21 spectrum was used for the calibration of AT2017gax from 2018-08-29.

2017-09-13:
Hip021545 GB spectra were not obtained with a proper AB offset and was not used in the reductions of AT2017gax. During the same night Hip029337 was observed for AT2017fzw and was used in the reductions of both science targets.

2017-09-13:
AT2017fzw sequence of 4 frames has 2 low quality frames in the end of the set. These were excluded from the reduction.

2017-09-29:
AT2017fzw sequence of 12 frames has 4 very low quality frames in the beginning of the set. These were excluded from the reduction.

2017-10-19:
Feige110 BG frames taken in the beginning of the night are not useful. However, observers obtained spectra of Feige110 again during the night and this data was used to reduce the science observations.

2017-10-19:
AT2017gax sequence of 8 frames has the first AB pair severely affected by instrumental noise, and these 2 frames were excluded from the reduction. Nonetheless, the resulting spectrum has a very low signal-to-noise ratio.

2017-10-29:
The PESSTO pipeline produces *_sc.fits files of HIP009911 in GB and GR, even if these are not science observations but a telluric standard.
HIP009911_20171019_GB_merge_58760_1_sc.fits
HIP009911_20171019_GR_merge_58760_1_sc.fits

2017-11-12:
AT2017fzw sequence of 12 frames has a very low signal-to-noise ratio with the target completely disappeared in the last 8 frames. These frames were excluded from the reduction.

2018-12-13:
SN2017iuk sequence of 6 frames has a very low signal-to-noise ratio in the last 2 frames. These frames were excluded from the reduction.

2018-02-04:
SN2018ec sequence of 12 frames has a major detector problem in the last 6 frames. These frames were excluded from the reduction.

2018-02-04:
SN2018oh sequence of 8 frames has very low signal-to-noise ratio in the last 6 frames. The observers note in the nightlog that thick clouds appeared, but the OB was not aborted. These 6 frames were excluded from the reduction.

2018-02-14:
SN2018ec sequence of 12 frames has a major detector problem in the top right quadrants. Frames B in the ABBA cycle were excluded from the reduction.

2018-02-14:
SN2018oh BG observations include the first sequence without signal, as noted by the observers in the nightlog. The following files were excluded and only the second sequence was reduced.
SOFI.2018-02-15T04:14:50.428.fits
SOFI.2018-02-15T04:20:42.440.fits
SOFI.2018-02-15T04:26:24.616.fits
SOFI.2018-02-15T04:32:13.348.fits

2018-03-25:
ASASSN-18fv GB spectrum might have an issue with the relative flux calibration?

2018-04-20:
LTT3218 spectroscopic standard star GB and GR observations failed. Not used in the reductions to calibrate observations of ASASSN-18fv, but instead observations of LTT3218 from 2018-04-07 were used for the flux calibration.

2018-05-13:
EG274 spectroscopic standard star GB and GR observations in AB mode obtained without dithering. Not used in the reductions to calibrate observations of ASASSN-18fv, but instead observations of LTT3218 from 2018-04-07 were used for the flux calibration.

2018-05-13:
ASASSN-18fv GB sequence had corrupted raw file in the tarball. Re-downloaded the following file from the ESO archive for the reductions:
SOFI.2018-05-13T05:03:13.702.fits

2018-08-11:
SN2018eov GB spectrum obtained with an ABBA dither that places the B position close to a position that is close to a nearby star in the A position. The B position frames were not included in the reduction.

2018-08-11:
Hip078425 telluric standard obtained to calibrate SN2018eov has a peculiar extracted GB spectral shape. SN2018eov calibrated instead with HIP071559 obtained to calibrate SN2018enc

2018-09-02:
The PESSTO pipeline produces *_sc.fits files of HIP3820 in GB, even if these are not science observations but a telluric standard.
HIP3820_20180902_GB_merge_58779_1_sc.fits

2018-09-02:
SN2018exc GB 2D frames have no signal at all. Either heavy clouds or the target is not on the slit. The first frames of the sequence suffer also from the bad top right quadrant. A *sc.fits file was somewhat arbitrarily produced, but should not be included in the data release?

2018-10-19:
SN2018hgc GB 2D frames have basically no signal at all. Either heavy clouds or the target is not on the slit. The last 6 frames of the sequence of 12 have some signal which was extracted. Data very likely not scientifically useful.

2019-01-16:
SN2018jky GB observations had initially detector noise problems which was noted also in the nightlog. Only frames obtained after the problem was fixed were used in the reductions. Affected frames are:
SOFI.2019-01-17T00:59:17.693.fits
SOFI.2019-01-17T01:06:29.023.fits

2019-01-16:
SN2018jky and SN2018hfm were observed with GB, however, no telluric or spectrophotometric standard star observations were obtained. GD71 and HIP48693 from 2018-11-17 were used to carry out the reductions.



Notes on the most prominent noise spikes in the reduced spectra:

2018-03-24:
A prominent noise spike in the telluric contamination region of the GB spectrum of ASASSN-18fv.

2017-08-29:
A prominent noise spike around 9793 Å in the GB spectrum of AT2017fzw.

2017-10-29:
A prominent noise spike around 12596 Å in the GB spectrum of AT2017fzw.

2017-11-09:
Prominent noise spikes around the telluric contamination region of the GB spectrum of AT2017fzw.

2017-08-19:
A prominent noise spike around 12603 Å in the GB spectrum of AT2017gah.

2017-12-13:
A prominent noise spike around 12605 Å in the GB spectrum of AT2017hcc.

2018-02-04:
Prominent noise spikes around the telluric contamination region of the GB spectrum of SN2018ec.

2018-02-14:
Prominent noise spikes around the telluric contamination region of the GB spectrum of SN2018ec.

2018-08-11:
A prominent noise spike around the telluric contamination region of the GB spectrum of SN2018eov.

2018-08-18:
A prominent noise spike around the telluric contamination region of the GB spectrum of SN2018exb.

2018-02-05:
Prominent noise spikes around the telluric contamination region of the GB spectrum of SN2018oh.

2018-02-06:
Prominent noise spikes around the telluric contamination region of the GB spectrum of SN2018oh.

2017-11-09:
A prominent noise spike at the red end of the GR spectrum of AT2017hcc.

2017-12-13:
A prominent noise spike at the red end of the GR spectrum of AT2017hcc.






Comments